OpenStudy (anonymous):

I need help with how to come up with 4 different fractions equal to 1

OpenStudy (anonymous):

Four fractions that add together to equal one?

OpenStudy (anonymous):

or 4 different fractions, each equal to 1 ? 2/2 3/3 4/4 5/5

OpenStudy (anonymous):

omgosh thank u so much

OpenStudy (anonymous):

o wait no four different fractions that r equal to one

OpenStudy (anonymous):

6/5 + 4/5 -3/4 -2/8

OpenStudy (anonymous):

or, the sum of the fractions below will be equal to 1: 1/10 + 2/10 + 3/10 + 4/10 = 10/10 = 1